Julio has four children whose ages are consecutive odd integers. If x represents the youngest child's age, what is the sum of his children ages
Answer by JulietG(1812)

A = x+6
B = x+4
C = x+2 (consecutive odd)
D = x (given)
A+B+C+D = sum of children's ages
The sum is 4x+12
What are the possible answers? We know that x is odd, so:
1,3,5,7 = 16 (x, the youngest is 1)
3,5,7,9 = 24
5,7,9,11 = 32
7,9,11,13 = 40
9,11,13,15 = 48
11,13,15,17 = 56
etc